OpenStudy (anonymous):

FAN MEDAL INSTAGRAM Given the system of equations presented here: 2x + 4y = 14 4x + y = 20 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (calculusxy):

we know that when a number and its opposite is added, then the result is 0. let's try to cancel x. we have 2x and 4x.

OpenStudy (anonymous):

Multiply both sides of the second equation by -4.

OpenStudy (calculusxy):

if we have -4 + 4 =0

OpenStudy (calculusxy):

thus multiply the first equation by -2 and the second one by -1

OpenStudy (anonymous):

Well these are the answers Multiply the second equation by −4 to get −16x − 4y = −80 Multiply the second equation by −1 to get −4x − y = −20 Multiply the first equation by 2 to get 4x + 8y = 28 Multiply the first equation by −1 to get −2x − 4y = −14

OpenStudy (anonymous):

x and y are fractions. {x = 33/7, y = 8/7}

OpenStudy (anonymous):

so which one would it be @robtobey

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

x = 33/7 y = 8/7 Plug the above values into each equation in order to verify their values.

OpenStudy (anonymous):

If each side of 2x + 4y = 14 is divided by 2, then x + 2y = 7 . Now the second equation can be multiplied by -2.